# The Atacama Cosmology Telescope DR6 and DESI: Structure formation over cosmic time with a measurement of the cross-correlation of CMB Lensing and Luminous Red Galaxies

### v1

@@ 1 Introduction
The matter distribution is typically characterized in terms of $\sigma_{8}$, the amplitude of matter density fluctuations smoothed on a scale of $8\,\text{Mpc}/h$. Weak lensing observables, in particular, measure degenerate combinations with the average matter density of the universe $\Omega_{m}$, e.g., $S_{8}=\sigma_{8}\sqrt{\Omega_{m}/0.3}$. Early observations of galaxy lensing with the CFHTLens survey [21] began to hint at a possible disagreement of this quantity between direct late-time observables and the primary CMB prediction [22, 23, 24]. Today, primary CMB measurements provide strong constraints on $S_{8}$ (derived through extrapolation to late times and assuming the $\Lambda$CDM model), e.g., $S_{8}=0.834\pm 0.016$ from Planck 2018 (PR3) [1], $S_{8}=0.827\pm 0.013$ from Planck NPIPE (PR4) [2], $S_{8}=0.830\pm 0.043$ from ACT DR4 [3], and $S_{8}=0.797\pm 0.042$ from the South Pole Telescope (SPT-3G, [4]) while measurements of the combination of galaxy weak lensing and galaxy clustering from surveys such as the Dark Energy Survey (DES, [11, 12]), the Kilo-Degree Survey (KiDS, [13, 14]), and the Hyper Suprime-Cam (HSC, [15, 16, 17, 18]) typically tend to find lower values, $S_{8}=0.776\pm 0.017$, $S_{8}=0.765^{+0.017}_{-0.016}$, and $S_{8}=0.775^{+0.043}_{-0.038}$ respectively. Low inferences are also found in full-shape analyses of galaxy clustering from the Baryon Oscillation Spectroscopic Survey (BOSS, e.g., [7, 25]), but clustering from the full Sloan Digital Sky Survey (SDSS, [26]) that includes BOSS data as well as the joint reanalysis of galaxy weak lensing data from DES Y3 and KiDS-1000 [27] find slightly higher values. Intriguingly, measurements of the CMB lensing power spectrum that best infer properties of structure at lower to intermediate redshifts $50.5<z<$ 5 [28] are in good agreement with the primary CMB: $S_{8}=0.831\pm 0.029$ from Planck PR4  [^1] The value of $S_{8}$ from Planck PR4 lensing was not explicitly provided in [29] but rather inferred from the chains provided in Section IV: https://github.com/carronj/planck_PR4_lensing [29], $S_{8}=0.840\pm 0.028$ from ACT DR6 [28, 30] and $S_{8}=0.836\pm 0.039$ from SPT-3G [31]. Galaxy cluster abundance measured by SPT ([10]) gives an intermediate value of $S_{8}=0.795\pm 0.029$, while an analysis using the first eROSITA All-Sky Survey (eRASS1, [9]) presents a higher value of $S_{8}=0.86\pm 0.01$.
@@ 2.1 DESI Luminous Red Galaxy sample
Figure 2: Left: The redshift distribution $dN(z)/dz$ of the DESI LRG galaxy redshift bins with the CMB lensing kernel shown in gray, showing ample overlap in redshifts between the two sets of cosmological probes. Right: Planck CMB prediction for the lensing power spectrum plotted against the lensing noise spectra of Planck PR4 (shown in blue) and ACT DR6 (shown in red). The lightly shaded bars in colors represent the fractional contribution to the cross-correlation $C_{L}^{\kappa g}$ signal-to-noise using covariances for Planck PR4 (blue) and ACT DR6 (red) and the same fiducial theory for both (see Appendix B for more details), showing us that Planck holds more constraining power than ACT until $L\approx 400$. The shaded bars in gray show angular multipoles excluded due to scale cuts chosen for the analysis (where the light gray band labeled “PR4 only” denotes an $L$ band used only for the Planck PR4 cross-correlation). Both: The colored bars and contours for both figures in addition to the gray CMB lensing kernel in the left figure are scaled to some arbitrary normalization factor for ease of visualization.
@@ 2.2.2 Planck PR4
Figure 3: The ACT DR6 lensing x DESI LRG cross-correlation angular power spectra and residuals, for all four redshift bins, with the diagonal elements of their simulation-based covariances used for their respective error bars. The Planck PR4 x DESI LRG cross-correlation spectra are shown as lighter-shaded bandpowers that are slightly shifted to the right from the ACT bandpowers for visual purposes. The signal-to-noise (SNR) ratio for each redshift bin is computed over the analysis $L$ range up to $L_{\rm max}=600$. The solid black curve in each plot is the power spectrum computed from the fiducial model using baseline best-fit cosmological parameters jointly fit to all four redshift bins, their auto-spectra, and their cross-correlations with ACT and Planck, within their respective analysis $L$ ranges. The best-fit spectra fit to 66 total degrees of freedom (d.o.f) results in a $\chi^{2}=54.1$ (15 d.o.f for $\chi^{2}=$ 11.5, 9.86, 16.1, 12.8 for each redshift bin fit independently). Assuming each free parameter removes exactly one degree of freedom, this leads to a probability-to-exceed (PTE) of 85.2%, demonstrating a good fit; [51] discusses the violation of this assumption for the case of prior-dominated parameters and provides a model fit PTE calculation.
@@ 3 CMB lensing tomography measurement
In spherical harmonic space, we perform an analysis of the two-point cross-correlation between the CMB lensing and the LRG overdensity fields as well as the two-point auto-correlation of the LRGs themselves. The technique to use varying redshift slices of galaxies in computing these two correlations jointly to constrain cosmology and the evolution of structure is known as CMB lensing tomography [69]. In this section, we describe the formalism for measuring the angular power spectra and its implementation. We use this implementation to measure power spectra for our data products as well as simulations which we use to estimate a transfer function and the data covariance.
@@
Figure 4: The DESI LRG angular auto power spectrum, with all four redshift bins and their simulation-based covariances used for their respective error bars. A fiducial value of the shot noise level estimated using a HEFT best-fit is subtracted for all four redshift bins, and is shown as colored dashed lines for the respective redshift bin. The power spectrum computed from the model described in the caption of Figure 3 (once again, fitting only to data in the non-gray regions) is shown in black; as demonstrated by the $\chi^{2}$ computation in Figure 3 ($\chi^{2}=54.1$, PTE = 85.2%) this is indeed a good fit.
@@ 3.1 Angular power spectrum
To prepare an $L$-dependent function (such as a theory spectrum) $C^{\prime}_{L}$ to compare directly with our estimation of the unbiased, binned angular power spectrum $C_{L_{b}}$, we convolve $C^{\prime}_{L}$ with our bandpower window functions, which applies the coupling, binning, and decoupling steps altogether; this procedure can be different from naively binning $C^{\prime}_{L}$ as the bandpower window functions correct for piecewise constant bins. This exact procedure is used to evaluate the likelihood for our analysis to compare our binned angular power spectrum data vector with a $C^{\prime}_{L}$ prediction from our theory model. For all purposes in this paper, the true angular power spectrum is computed by using the compute_full_master method in NaMaster that implements this pseudo-power spectrum estimator.
@@ 5.2 Theory model
HEFT allows us to parameterize cosmological power spectra as a linear combination of the CDM + baryon power spectrum $P_{cb}(k)$ and various intermediate component-basis spectra $P_{i,j}(k)$ that capture two-point correlations between different overdensities expressed in the Lagrangian bias formalism. Predictions for the intermediate power spectra are computed efficiently by an emulator trained on the Aemulus-$\nu$ simulations [98], and then Limber integrated over the line-of-sight to obtain predictions for the observables $C_{L}^{gg}$ and $C_{L}^{\kappa g}$. As the theory power spectrum depends linearly on the counterterms for the galaxy auto-spectrum and the cross-spectrum with matter as well as the shot noise $SN$, we can assume a Gaussian prior for these linear parameters and analytically marginalize our likelihood with respect to them. Further details of the marginalization procedure, and its implementation in our likelihood can be seen in [51].
@@ 5.3 Cosmological parameterization and priors
Table 2: Parameters and priors used in this work and [51]. Uniform priors from $x_{1}$ to $x_{2}$ are denoted with $\mathcal{U}(x_{1},x_{2})$ and Gaussian priors with mean $\mu$ and standard deviation $\sigma$ are shown as $\mathcal{N}(\mu,\sigma)$. Nuisance parameters $b_{1},b_{2},b_{s}$ are all bias parameters for the HEFT theory model; counterterms are represented with $\alpha_{0}$ and $\epsilon$; and $\mu_{i}$ is the magnification bias for galaxy redshift bin $z_{i}$. Only the shot noise spectrum $N_{L}^{gg}$ and magnification bias $\mu_{i}$ have redshift bin-dependent priors, with $\mu$ and $\sigma$ shown respectively for bins 1, 2, 3, and 4.
@@
For the HEFT model, we put priors on analytically marginalized parameters ($\alpha_{a}$ for the auto counterterm, $\epsilon$ as a parameterization of the cross counterterm, and $N_{L}^{gg}$ for the shot noise), the Lagrangian bias parameters $b_{1}$, $b_{2}$, and $b_{s}$ (up to second or 1-loop order), and the magnification bias $\mu$. These are discussed in further detail in [51].
@@ 5.4 Parameter inference
$-2\ln\mathcal{L}\propto\begin{bmatrix}\hat{C}_{L}^{\kappa g}-C_{L}^{\kappa g}(\bm{\theta})\\ \hat{C}_{L}^{gg}-C_{L}^{gg}(\bm{\theta})\\ \end{bmatrix}^{T}\begin{bmatrix}\text{Cov}(C_{L}^{\kappa g},C_{L^{\prime}}^{\kappa g})&\text{Cov}(C_{L}^{\kappa g},C_{L^{\prime}}^{gg})\\ \text{Cov}(C_{L}^{gg},C_{L^{\prime}}^{\kappa g})&\text{Cov}(C_{L}^{gg},C_{L^{\prime}}^{gg})\\ \end{bmatrix}\begin{bmatrix}\hat{C}_{L^{\prime}}^{\kappa g}-C_{L^{\prime}}^{\kappa g}(\bm{\theta})\\ \hat{C}_{L^{\prime}}^{gg}-C_{L^{\prime}}^{gg}(\bm{\theta})\\ \end{bmatrix}$ (5.1)
@@ 5.6 Results
We show $S_{8}^{\times}(z)$ for each of our 4 redshift bins in Figure 9 by using the growth function of the Planck PR3 cosmology to rescale the ACT + Planck $S_{8}^{\times}$ means and errors from redshift zero to the effective redshifts (see Table 1 in [51]) for each bin (angle bracketed values represent the Planck PR4 [2] prediction of $S_{8}^{\times}(z)$):

### v2

@@ 1 Introduction
The matter distribution is typically characterized in terms of $\sigma_{8}$, the amplitude of matter density fluctuations smoothed on a scale of $8\,\text{Mpc}/h$. Weak lensing observables, in particular, measure degenerate combinations with the average matter density of the universe $\Omega_{m}$, e.g., $S_{8}=\sigma_{8}\sqrt{\Omega_{m}/0.3}$. Early observations of galaxy lensing with the CFHTLens survey [21] began to hint at a possible disagreement of this quantity between direct late-time observables and the primary CMB prediction [22, 23, 24]. Today, primary CMB measurements provide strong constraints on $S_{8}$ (derived through extrapolation to late times and assuming the $\Lambda$CDM model), e.g., $S_{8}=0.834\pm 0.016$ from Planck 2018 (PR3) [1], $S_{8}=0.827\pm 0.013$ from Planck NPIPE (PR4) [2], $S_{8}=0.830\pm 0.043$ from ACT DR4 [3], and $S_{8}=0.797\pm 0.042$ from the South Pole Telescope (SPT-3G, [4]) while measurements of the combination of galaxy weak lensing and galaxy clustering from surveys such as the Dark Energy Survey (DES, [11, 12]), the Kilo-Degree Survey (KiDS, [13, 14]), and the Hyper Suprime-Cam (HSC, [15, 16, 17, 18]) typically tend to find lower values, $S_{8}=0.776\pm 0.017$, $S_{8}=0.765^{+0.017}_{-0.016}$, and $S_{8}=0.775^{+0.043}_{-0.038}$ respectively. Low inferences are also found in full-shape analyses of galaxy clustering from the Baryon Oscillation Spectroscopic Survey (BOSS, e.g., [7, 25]), but clustering from the full Sloan Digital Sky Survey (SDSS, [26]) that includes BOSS data as well as the joint reanalysis of galaxy weak lensing data from DES Y3 and KiDS-1000 [27] find slightly higher values. Intriguingly, measurements of the CMB lensing power spectrum that best infer properties of structure at intermediate redshifts $50.5<z<$ 5 [28] are in good agreement with the primary CMB: $S_{8}=0.831\pm 0.029$ from Planck PR4  [^1] The value of $S_{8}$ from Planck PR4 lensing was not explicitly provided in [29] but rather inferred from the chains provided in Section IV: https://github.com/carronj/planck_PR4_lensing [29], $S_{8}=0.840\pm 0.028$ from ACT DR6 [28, 30] and $S_{8}=0.836\pm 0.039$ from SPT-3G [31]. Galaxy cluster abundance measured by SPT ([10]) gives an intermediate value of $S_{8}=0.795\pm 0.029$, while an analysis using the first eROSITA All-Sky Survey (eRASS1, [9]) presents a higher value of $S_{8}=0.86\pm 0.01$.
@@ 2.1 DESI Luminous Red Galaxy sample
Figure 2: Left: The redshift distribution $dN(z)/dz$ of the DESI LRG galaxy redshift bins with the CMB lensing kernel shown in gray, showing ample overlap in redshifts between the two sets of cosmological probes. Right: Planck CMB prediction for the lensing power spectrum plotted against the lensing noise spectra of Planck PR4 (shown in blue) and ACT DR6 (shown in red). The lightly shaded bars in colors represent the fractional contribution to the cross-correlation $C_{L}^{\kappa g}$ signal-to-noise using covariances for Planck PR4 (blue) and ACT DR6 (red) and the same fiducial theory for both (see Appendix B for more details), showing us that Planck holds more constraining power than ACT until $L\approx 400$. The shaded bars in gray show angular multipoles excluded due to scale cuts chosen for the analysis (where the light gray band labeled “PR4 only” denotes an $L$ band included only for the Planck PR4 cross-correlation). Both: The colored bars and contours for both figures in addition to the gray CMB lensing kernel in the left figure are scaled to some arbitrary normalization factor for ease of visualization.
@@ 2.2.2 Planck PR4
Figure 3: The ACT DR6 lensing x DESI LRG cross-correlation angular power spectra and residuals, for all four redshift bins, with the diagonal elements of their simulation-based covariances used for their respective error bars. The Planck PR4 x DESI LRG cross-correlation spectra are shown as lighter-shaded bandpowers that are slightly shifted to the right from the ACT bandpowers for visual purposes. The signal-to-noise (SNR) ratio for each redshift bin is computed over the analysis $L$ range up to $L_{\rm max}=600$. The solid black curve in each plot is the power spectrum computed from the fiducial model using baseline best-fit cosmological parameters jointly fit to all four redshift bins, their auto-spectra, and their cross-correlations with ACT and Planck, within their respective analysis $L$ ranges. The best-fit spectra fit to 66 total degrees of freedom (computed from subtracting the number of free parameters of the model fit from the total number of bandpowers being fit to, henceforth “d.o.f”) results in a $\chi^{2}=54.1$ (15 d.o.f for $\chi^{2}=$ 11.5, 9.86, 16.1, 12.8 for each redshift bin fit independently). Assuming each free parameter removes exactly one degree of freedom, this leads to a probability-to-exceed (PTE) of 85.2%, demonstrating a good fit; [51] discusses the violation of this assumption for the case of prior-dominated parameters and provides a model fit PTE calculation.
@@ 3 CMB lensing tomography measurement
In spherical harmonic space, we perform an analysis of the two-point cross-correlation between the CMB lensing and the LRG overdensity fields as well as the two-point auto-correlation of the LRGs themselves. To constrain cosmology and the evolution of structure, we use a technique to use varying redshift slices of galaxies in computing these two correlations jointly known as CMB lensing tomography [69]. In this section, we describe the formalism for measuring the angular power spectra and its implementation. We use this implementation to measure power spectra for our data products as well as simulations which we use to estimate a transfer function and the data covariance.
@@
Figure 4: The DESI LRG angular auto power spectrum, with all four redshift bins and the diagonals of their simulation-based covariances used for their respective error bars. A fiducial value of the shot noise level estimated using a HEFT best-fit is subtracted for all four redshift bins, and is shown as colored dashed lines for the respective redshift bin. The power spectrum computed from the model described in the caption of Figure 3 (once again, fitting only to data in the non-gray regions) is shown in black; as demonstrated by the $\chi^{2}$ computation in Figure 3 ($\chi^{2}=54.1$, PTE = 85.2%) this is indeed a good fit.
@@ 3.1 Angular power spectrum
To prepare an $L$-dependent function (such as a theory spectrum) $C^{\prime}_{L}$ to compare directly with our estimation of the unbiased, binned angular power spectrum $C_{L_{b}}$, we convolve $C^{\prime}_{L}$ with our bandpower window functions, which applies the coupling, binning, and decoupling steps altogether; this procedure can be different from naively binning $C^{\prime}_{L}$ as the bandpower window functions correct for piecewise constant bins. The same procedure is used to evaluate the likelihood for our analysis to compare our binned angular power spectrum data vector with a $C^{\prime}_{L}$ prediction from our theory model. For all purposes in this paper, the true angular power spectrum is computed by using the compute_full_master method in NaMaster that implements this pseudo-power spectrum estimator.
@@ 5.2 Theory model
HEFT allows us to parameterize cosmological power spectra as a linear combination of the CDM + baryon power spectrum $P_{cb}(k)$ and various intermediate component-basis spectra $P_{i,j}(k)$ that capture two-point correlations between different overdensities expressed in the Lagrangian bias formalism. To 1-loop or second order, this linear combination is expressed using a set of Lagrangian bias parameters $b_{i}$ for $i=1,2,s$ that quantify the contribution of the CDM + baryon overdensity fields $\delta_{cb}$, $\delta_{cb}^{2}$, and the tidal shear field $s_{cb}$ respectively. As highlighted in [51], we also use counterterms $\alpha$ to capture interactions with the derivative field $\nabla^{2}\delta_{cb}$ and other small-scale stochastic components. Using these bias parameters that are independently defined and varied per redshift bin along with cosmological parameters as inputs, predictions for the intermediate power spectra are computed efficiently by an emulator trained on the Aemulus-$\nu$ simulations [98] which model, and then Limber integrated over the line-of-sight to obtain predictions for the observables $C_{L}^{gg}$ and $C_{L}^{\kappa g}$. As the theory power spectrum depends linearly on the counterterms for the galaxy auto-spectrum and the cross-spectrum with matter as well as the shot noise $SN$, we can assume a Gaussian prior for these linear parameters and analytically marginalize our likelihood with respect to them. Further details of the marginalization procedure, and its implementation in our likelihood can be seen in [51].
@@ 5.3 Cosmological parameterization and priors
Table 2: Parameters and priors used in this work and [51]. Uniform priors from $x_{1}$ to $x_{2}$ are denoted with $\mathcal{U}(x_{1},x_{2})$ and Gaussian priors with mean $\mu$ and standard deviation $\sigma$ are shown as $\mathcal{N}(\mu,\sigma)$. Nuisance parameters $b_{1},b_{2},b_{s}$ are all bias parameters for the HEFT theory model; counterterms are represented with $\alpha_{a}$ and $\epsilon$; and $\mu_{i}$ is the magnification bias for galaxy redshift bin $z_{i}$. Only the shot noise spectrum $N_{L}^{gg}$ and magnification bias $\mu_{i}$ have redshift bin-dependent priors, with $\mu$ and $\sigma$ shown respectively for bins 1, 2, 3, and 4.
@@
For the HEFT model, we put priors on analytically marginalized parameters ($\alpha_{a}$ for the auto counterterm, $\epsilon$ as a parameterization of the cross counterterm, and $N_{L}^{gg}$ for the shot noise), the Lagrangian bias parameters $b_{1}$, $b_{2}$, and $b_{s}$ (up to second or 1-loop order), and the magnification bias $\mu$. We put relatively uninformative priors on all of these HEFT parameters except for $b_{s}$ and $\epsilon$, where the former is found to share a strong degeneracy with $b_{2}$ and the latter is chosen to appropriately represent the size of small-scale effects we expect from baryonic feedback [32] and our usage of the Aemulus-$\nu$ simulations. These are discussed in further detail in [51].
@@ 5.4 Parameter inference
$-2\ln\mathcal{L}\propto\begin{bmatrix}\hat{C}_{L}^{\kappa g}-C_{L}^{\kappa g}(\bm{\theta})\\ \hat{C}_{L}^{gg}-C_{L}^{gg}(\bm{\theta})\\ \end{bmatrix}^{T}\begin{bmatrix}\text{Cov}(C_{L}^{\kappa g},C_{L^{\prime}}^{\kappa g})&\text{Cov}(C_{L}^{\kappa g},C_{L^{\prime}}^{gg})\\ \text{Cov}(C_{L}^{gg},C_{L^{\prime}}^{\kappa g})&\text{Cov}(C_{L}^{gg},C_{L^{\prime}}^{gg})\\ \end{bmatrix}^{-1}\begin{bmatrix}\hat{C}_{L^{\prime}}^{\kappa g}-C_{L^{\prime}}^{\kappa g}(\bm{\theta})\\ \hat{C}_{L^{\prime}}^{gg}-C_{L^{\prime}}^{gg}(\bm{\theta})\\ \end{bmatrix}$ (5.1)
@@ 5.6 Results
We show $S_{8}^{\times}(z)$ for each of our 4 redshift bins in Figure 9 by rescaling the ACT + Planck $S_{8}^{\times}$ means and errors from redshift zero to their effective redshifts (see Table 1 in [51]); curly bracketed values represent the Planck PR4 [2] primary CMB prediction of $S_{8}^{\times}(z)$: